# Distilling Calibration via Conformalized Credal Inference

### v2

@@ Abstract
Index Terms:
Calibration, distillation, conformal prediction, credal inference
{strip}
Figure 1: Given an input $x$, the predictive distribution ideally coincides with that of a large-scale cloud-based model $p^{*}(\cdot|x)$. In the setting studied in this work, a small-scale edge-based model produces a probabilistic distribution $p(\cdot|x)$ that deviates from the reference distribution $p^{*}(\cdot|x)$, and is thus uncalibrated. The proposed conformalized credal inference-based scheme post-processes the small-scale edge model output $p(\cdot|x)$ via a simple thresholding mechanism to produce a subset $\Gamma(x)$ in the simplex of predictive distributions, with the guarantee of containing the reference distribution $p^{*}(\cdot|x)$ with probability $1-\epsilon$. A final calibrated predictive distribution can be obtained via ensembling or via other combining mechanisms.
@@ III-A Credal Sets
where $f(\cdot)$ is a convex function satisfying the properties (i) $f(1)=0$, and (ii) $00\cdot f(0/0)=$ 0 [58, 59]. The class of $f$-divergences encompasses a variety of divergence measures, including the Kullback–Leibler (KL) divergence and the Tsallis divergence.
@@ III-C Theoretical Reliability Guarantees
$\Pr\left[p^{*}(\cdot|x)\in\Gamma(x)\right]\geq 1-\epsilon,\text{ for any test input }x.$ (8)
@@ IV-A Intersection Probability
The distribution in (10) ensures that each class $y$ is treated equally. This is done by choosing the probability $q^{*}(y|x)$ to be within the interval $[q_{L_{y}},q_{U_{y}}]$ at the same relative position determined by the fraction $\beta$ for all $y\in\mathcal{K}$.
@@ IV-B Low-Complexity Bayesian Learning via the Laplace Approximation
To elaborate, assume a parametric small-scale model
$p(y|x)=p(y|x,\hat{\theta}),$ (13)
@@ V-C Image Classification
The inefficiency and coverage of CD-CI as a function of the test accuracy of the small-scale model are shown in Fig. 3. Note that the accuracy of the small-scale model ranges from $70\%$ to $90\%$, while the large-scale model obtains accuracy $95\%$. We set the target miscoverage rate as $\epsilon=0.1$. Improving the accuracy of the small-scale model is seen to be instrumental in enhancing the efficiency of the conformalized credal set. Furthermore, CD-CI maintains a coverage rate close to the target level $0.91-\epsilon=$ 0.9, for all small-scale models, validating Theorem 1.
[…]
@@
[…]
Fig. 4 and Fig. 5 also illustrate the advantages of constraining the credal set through the choice of a value of $\alpha$ larger than $\alpha=1$, thus moving beyond the KL divergence. To investigate the impact of the divergence parameter $\alpha$, the ECE and the accuracy are shown in Fig. 6 and Fig. 7, respectively, as a function of $\alpha$ for $0.91-\epsilon=$ 0.9. In these figures, we evaluate and compare three approaches for deriving a single predictive distribution, namely (i) the intersection probability (10); (ii) the ensemble distribution $\mathbb{E}_{\Gamma(x)}[q(\cdot|x)]$, where the average is over a uniform distribution in the credal set $\Gamma(x)$; and (iii) the maximum Shannon entropy distribution $\max_{q(\cdot|x)\in\Gamma(x)}[H(q(\cdot|x)]$ [25].
@@ V-D Natural Language Classification
Firstly, to visualize the calibration benefits of CD-CI, Fig. 8 plots the confidence histograms evaluated over test data $(x,y)$ for the large-scale model predictive distribution $p^{*}(y|x)$, the small-scale model $p(y|x)$, and CD-CI $q^{*}(y|x)$ on the SNLI data set, using $\alpha=1$ with target coverage rate $0.91-\epsilon=$ 0.9. As illustrated, while the small-scale model is overconfident, CD-CI successfully recalibrate the model, ensuring decisions that are calibrated to a level similar to the reference large-scale model.
Figure 8: Confidence histograms for the large-scale model predictive distribution $p^{*}(y|x)$, for the small-scale model $p(y|x)$, and for CD-CI $q^{*}(y|x)$ evaluated on SNLI data set, using $\alpha=1$ with target coverage rate $0.91-\epsilon=$ 0.9.
@@
[…]
Figure 10: ECE versus target coverage rate $1-\epsilon$ for different values of the $\alpha$ for $\alpha$-divergence used in (4) on the SNLI data set. The dashed lines report the ECE performance of the large-scale model predictive distribution $p^{*}(\cdot|x)$, and of the small-scale model $p(\cdot|x)$.
@@
Fig. 10 shows the ECE of CD-CI as a function of the target coverage level $1-\epsilon$, ranging from $80\%$ to $95\%$, for different values of $\alpha$. These results confirm the general conclusions from Fig. 4. In fact, in this experiment, the benefits of CD-CI are seen to be even more robust to the choice of hyperparameters $1-\epsilon$ and $\alpha$ in terms of the ECE. Note that the Laplace approximation was infeasible for this task, given the adoption of a Transformer network architecture. Overall, CD-CI achieves an improvement of approximately $5\%$ in ECE compared to the original small-scale model.

### v4

@@ III-A Credal Sets
where $f(\cdot)$ is a convex function satisfying the properties (i) $f(1)=0$, and (ii) $0\cdot f(0/0)=0$ [58, 59]. The class of $f$-divergences encompasses a variety of divergence measures, including the Kullback–Leibler (KL) divergence and the Tsallis divergence.
@@ III-C Theoretical Reliability Guarantees
$\Pr\left[p^{*}(\cdot|x)\in\Gamma(x)\right]\geq 1-\epsilon.$ (8)
@@ IV-A Intersection Probability
The distribution in (10) ensures that each class $y$ is treated equally. This is done by choosing the probability $q^{*}(y|x)$ to be within the interval $q_{L_{y}},q_{U_{y}}]$ at the same relative position determined by the fraction $\beta$ for all $y\in\mathcal{K}$.
@@ IV-B Low-Complexity Bayesian Learning via the Laplace Approximation
Figure 2: Test input $x$, reference distribution $p^{*}(\cdot|x)$ from the large-scale model, and credal sets produced by CD-CI for small-scale models with different accuracy on the CIFAR-10 data set with classes $\{\text{airplane, automobile, bird}\}$ using the KL divergence in (4) with target coverage rate $1-\epsilon=0.9$. Note that the large-model distribution $p^{*}(\cdot|x)$ is marked as red point in the simplex.
To elaborate, assume a parametric small-scale model
@@ V-C Image Classification
The inefficiency and coverage of CD-CI as a function of the test accuracy of the small-scale model are shown in Fig. 3. Note that the accuracy of the small-scale model ranges from $70\%$ to $90\%$, while the large-scale model obtains accuracy $95\%$. We set the target miscoverage rate as $\epsilon=0.1$. Improving the accuracy of the small-scale model is seen to be instrumental in enhancing the efficiency of the conformalized credal set. Furthermore, CD-CI maintains a coverage rate close to the target level $1-\epsilon=0.9$, for all small-scale models, validating Theorem 1.
Figure 3: Coverage and inefficiency versus the small-scale models accuracy on the CIFAR-10 data set with classes $\{\text{airplane, automobile, bird}\}$ using the KL divergence in (4) with target coverage rate $1-\epsilon=0.9$. The accuracy of the large-scale model, ResNet-18 network, is $95.58\%$, and the accuracy of the small-scale model, Mini-VGG-8, is controlled by training over different numbers of iterations.
@@
Figure 4: ECE versus target coverage rate $1-\epsilon$ for different values of $\alpha$ for the $\alpha$-divergence used in (4) on the CIFAR-10 data set with classes $\{\text{airplane, automobile, bird}\}$. The dashed lines report the ECE performance of the large-scale model predictive distribution $p^{*}(\cdot|x)$, of the small-scale model $p(\cdot|x)$, and of the Laplace approximation method $q^{\text{La}}(\cdot|x)$ in (15).
Figure 5: Accuracy versus target coverage rate $1-\epsilon$ for different values of $\alpha$ for the $\alpha$-divergence used in (4) on the CIFAR-10 data set with classes $\{\text{airplane, automobile, bird}\}$. The dashed lines report the accuracy performance of the large-scale model predictive distribution $p^{*}(\cdot|x)$, and of the small-scale model $p(\cdot|x)$. Note that there is no change to accuracy when applying Laplace approximation in a post-processing way.
Figure 6: ECE versus the values of $\alpha$ for the $\alpha$-divergence used in (4) for different target coverage rate $1-\epsilon$ on the CIFAR-10 data set with classes $\{\text{airplane, automobile, bird}\}$. The blue line represents the ECE performance of the predictive distribution $q^{*}(y|x)$ (10), in yellow the ECE of the ensemble distribution $\mathbb{E}_{\Gamma(x)}[q(\cdot|x)]$, and in cyan the ECE of the maximum Shannon entropy distribution $\max_{q(\cdot|x)\in\Gamma(x)}[H(q(\cdot|x)]$.
Figure 7: Accuracy versus the values of $\alpha$ for the $\alpha$-divergence used in (4) for different target coverage rate $1-\epsilon$ on the CIFAR-10 data set with classes $\{\text{airplane, automobile, bird}\}$. Legend is as in Figure 6.
Fig. 4 and Fig. 5 also illustrate the advantages of constraining the credal set through the choice of a value of $\alpha$ larger than $\alpha=1$, thus moving beyond the KL divergence. To investigate the impact of the divergence parameter $\alpha$, the ECE and the accuracy are shown in Fig. 6 and Fig. 7, respectively, as a function of $\alpha$ for $1-\epsilon=0.9$. In these figures, we evaluate and compare three approaches for deriving a single predictive distribution, namely (i) the intersection probability (10); (ii) the ensemble distribution $\mathbb{E}_{\Gamma(x)}[q(\cdot|x)]$, where the average is over a uniform distribution in the credal set $\Gamma(x)$; and (iii) the maximum Shannon entropy distribution $\max_{q(\cdot|x)\in\Gamma(x)}[H(q(\cdot|x)]$ [25].
@@ V-D Natural Language Classification
Firstly, to visualize the calibration benefits of CD-CI, Fig. 8 plots the confidence histograms evaluated over test data $(x,y)$ for the large-scale model predictive distribution $p^{*}(y|x)$, the small-scale model $p(y|x)$, and CD-CI $q^{*}(y|x)$ on the SNLI data set, using $\alpha=1$ with target coverage rate $1-\epsilon=0.9$. As illustrated, while the small-scale model is overconfident, CD-CI successfully recalibrate the model, ensuring decisions that are calibrated to a level similar to the reference large-scale model.
Figure 8: Confidence histograms for the large-scale model predictive distribution $p^{*}(y|x)$, for the small-scale model $p(y|x)$, and for CD-CI $q^{*}(y|x)$ evaluated on SNLI data set, using $\alpha=1$ with target coverage rate $1-\epsilon=0.9$.
@@
Figure 9: Coverage and inefficiency versus the small-scale models accuracy on the SNLI data set using the KL divergence in (4) with target coverage rate $1-\epsilon=0.9$. The accuracy of the large-scale model is $91.58\%$. Note that INTXX denotes the block-floating point format, in which each element of the block is expressed via INTXX precision (see e.g., [56, 68]).
Figure 10: ECE versus target coverage rate $1-\epsilon$ for different values of the $\alpha$ for $\alpha$-divergence used in (4) on the SNLI data set. The dashed lines report the ECE performance of the large-scale model predictive distribution $p^{*}(\cdot|x)$, of the small-scale model $p(\cdot|x)$, and of the Laplace approximation method $q^{\text{La}}(\cdot|x)$ in (15).
@@
Fig. 10 shows the ECE of CD-CI as a function of the target coverage level $1-\epsilon$, ranging from $80\%$ to $95\%$, for different values of $\alpha$. These results confirm the general conclusions from Fig. 4. In fact, in this experiment, the benefits of CD-CI are seen to be even more robust to the choice of hyperparameters $1-\epsilon$ and $\alpha$ in terms of the ECE. Overall, CD-CI achieves an improvement of approximately $5\%$ and $2\%$ in ECE compared to the original small-scale model and the Laplace approximation.